OpenStudy (anonymous):

olve the following equation: (4x – 7)(2x – 3) = 0 A. {-7/4, -3/2} B. {-4/7, -2/3} C. {7/4, 3/2} D. {4/7, 2/3}

OpenStudy (texaschic101):

4x - 7 = 0 4x = 7 x = 7/4 2x - 3 = 0 2x = 3 x = 3/2 answer C

OpenStudy (texaschic101):

you set each of them up equal to zero, then solve for x. It is really very simple. Do you have any questions because I will be happy to answer them.